# A $2$-distance $(2\Delta+7)$-coloring of planar graphs

### v2

@@ 2.2. The case $\Delta=7$
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=2$, then $v$ has either a $7(5^{-})$-neighbour or two $7(6)$-neighbours by Lemma 2.20(a). In both cases, $v$ receives at least $\frac{1}{4}=\min\{\frac{1}{4},2\times\frac{1}{6}\}$ from its $7$-neighbours by applying R6-R8. On the other hand, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{4}+\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=3$, then $v$ has two $7(6^{-})$-neighbours by Lemma 2.20(b). By applying R6-R8, $v$ receives totally at least $\frac{1}{3}$ from its $7$-neighbours. Thus, $\mu^{*}(v)\geq\frac{1}{3}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
@@
(3.2). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.23(d), $v$ has two neighbours different from $5^{-}$- and 6(6)666(6)6 ( 6 )-vertices, i.e., $v$ has two neighbours consisting of $6(5^{-})$- or $7$-vertices. It follows that $v$ receives totally at least $\frac{1}{6}=\min\{2\times\frac{1}{9},\frac{1}{9}+\frac{1}{12},2\times\frac{1}{12}\}$ from its $6^{+}$-neighbours by R3-R9, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+\frac{1}{6}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Next we suppose that $m_{4}(v)=1$. By Lemma 2.23(e), $v$ has at least three $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. If $v$ has no 7(7)777(7)7 ( 7 )-neighbour, then $v$ receives totally at least $3\times\frac{1}{9}$ from its $6^{+}$-neighbours by R3-R8. If $v$ has a 7(7)777(7)7 ( 7 )-neighbours, then $v$ has four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex by Lemma 2.23(f), and so $v$ receives totally at least $4\times\frac{1}{12}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.3.1). Let $n_{5}(v)=0$. By Lemma 2.24(b), $v$ has at most one 6(6)666(6)6 ( 6 )-neighbour, i.e., $v$ has at least four $6^{+}$-neighbours different from 6(6)666(6)6 ( 6 )-vertex. First, suppose that $v$ has a 6(6)666(6)6 ( 6 )-neighbour. If $v$ has no $6(5^{-})$-neighbour, then all the neighbours of $v$ different from 6(6)666(6)6 ( 6 )-vertex are $7$-vertices, and by applying R6-R9, $v$ receives totally at least $4\times\frac{1}{6}$ from its $7$-neighbours. So, $\mu^{*}(v)\geq 1+4\times\frac{1}{6}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has a $6(5^{-})$-neighbour, then $v$ has three $7$-neighbours, at least one of which is a $7(5^{-})$-vertex by Lemma 2.24(d). Then $v$ receives at least $\frac{1}{9}$ from its $6(5^{-})$-neighbour by R3-R5, at least $\frac{2}{9}$ from each of its $7(5^{-})$-neighbours by R6-R7, at least $\frac{1}{6}$ from each of its other $7$-neighbours by R6-R9. So, $v$ receives totally at least $\frac{1}{9}+\frac{2}{9}+2\times\frac{1}{6}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Suppose now that $v$ has no 6(6)666(6)6 ( 6 )-neighbour. That is, all neighbours of $v$ are $6^{+}$-vertices different from 6(6)666(6)6 ( 6 )-vertex. By Lemma 2.24(b), $v$ has at least one $7$-neighbour, say $x$. Note that if $v$ has exactly four $6$-neighbours, then $x$ must be a $7(5^{-})$-vertex by Lemma 2.24(c). This means that $v$ has either four $6$-neighbours and one $7(5^{-})$-neighbour or at most three $6$-neighbours and two $7$-neighbours. In each case, $v$ receives totally at least $\frac{2}{3}=\min\{4\times\frac{1}{9}+\frac{2}{9},3\times\frac{1}{9}+2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3.2). Let $n_{5}(v)=1$. Notice that $v$ has no 6(6)666(6)6 ( 6 )-neighbours by Lemma 2.24(e), i.e., all neighbours of $v$ but one are $7$- or $6(5^{-})$-vertices. Moreover, $v$ has at most two $6(5^{-})$-neighbours by Lemma 2.24(e). We distinguish three cases according to the number of $6(5^{-})$-vertices adjacent to $v$ as follows.
@@
Suppose next that $v$ has exactly one $6(5^{-})$-neighbour. Then $v$ has three $7$-neighbours, one of which is a $7(5^{-})$-vertex by Lemma 2.24(f). Thus, $v$ receives totally at least $\frac{1}{9}+2\times\frac{1}{6}+\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R9. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Finally, suppose that $v$ has exactly two $6(5^{-})$-neighbours. Then $v$ has two $7(5^{-})$-neighbours by Lemma 2.24(g). So, $v$ receives totally at least $2\times\frac{1}{9}+2\times\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R7. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(5.1). Let $m_{3}(v)=4$. By Lemma 2.26(b), $v$ has at most six $4$-neighbours. If $n_{4}(v)\leq 4$, then $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), and at most $\frac{2}{9}$ to each of its other neighbours by R7. Consequently, $\mu^{*}(v)\geq 3-4\times\frac{1}{4}-3\times\frac{2}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $5\leq n_{4}(v)\leq 6$. Clearly, $v$ has no 5(5)555(5)5 ( 5 )-neighbour, since a $4$ vertex is not adjacent to any 5(5)555(5)5 ( 5 )-vertex by Lemma 2.19(a). If $v$ has a $3$-neighbour $x$, then the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.17. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-faces by R2. Thus $\mu^{*}(v)\geq 3+\frac{1}{5}-6\times\frac{1}{4}-\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{5}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has no $3$-neighbour, then $\mu^{*}(v)\geq 3-6\times\frac{1}{4}-\frac{1}{7}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(5.2). Let $m_{3}(v)=5$. Notice that $v$ has at most one $3$-neighbour, since a $3$-vertex does not incident to any $3$-faces by Lemma 2.17. First we assume that $v$ has a $3$-neighbour $x$. In such a case, all $3$-faces incident to $f$ must have a consecutive ordering. Moreover, the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.17. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-faces by R2. Observe first that if $v$ has three $4$-neighbours, then at least one of them must be support vertex of $v$. In such a case, we have $d_{2}(v)\geq 21$ by Lemma 2.26(a). This implies that $v$ has at most four poor neighbours, and recall that a poor vertex receives at most $\frac{1}{4}$ from $v$ by R7(b),(c). It then follows that $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its poor neighbours by R7(b),(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e). We now suppose that $v$ has at most two $4$-neighbours. It then follows from Lemma 2.26(c) that $v$ has at most four poor neighbours. Thus, similarly as above, we have $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-2\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
@@
We may further assume that $v$ has no $3$-neighbours. By Lemma 2.26(b), $v$ has at most six $4$-neighbours, i.e., $n_{4}(v)\leq 6$. On the other hand, recall that $m_{3}(v)=5$. Therefore, if $v$ has more than four $4$-neighbours, then there exists $v_{i}\in N(v)$ such that $vv_{i}$ is contained in two $3$-faces, i.e., $v_{i}$ is a support vertex. It follows from Lemma 2.26(a) that we have $d_{2}(v)\geq 21$ when $v$ has a support vertex.
@@
(5.2.1). Let $n_{4}(v)\leq 2$. Note that $v$ has at most four poor neighbours by Lemma 2.26(c). Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{4}-2\times\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
@@
(5.2.2). Let $n_{4}(v)=3$. Obviously, $v$ has at most two 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.26(e). Suppose first that $v$ has exactly two 5(5)555(5)5 ( 5 )-neighbours $x,y$. Since a $4$-vertex has no 5(5)555(5)5 ( 5 )-neighbours by Lemma 2.19(a), we deduce that $x$ and $y$ are adjacent. On the other hand, a 5(5)555(5)5 ( 5 )-vertex has at most one bad $5$-neighbour by Lemma 2.24(b), i.e., a 5(5)555(5)5 ( 5 )-vertex cannot have both a 5(5)555(5)5 ( 5 )-neighbour and a $5(4)$-neighbour simultaneously. This implies that the neighbours of $v$ other than $4$- and 5(5)555(5)5 ( 5 )-vertices are $5(3^{-})$- or $6^{+}$-vertices. Thus we have $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). Suppose now that $v$ has at most one 5(5)555(5)5 ( 5 )-neighbour. In such a case, we have again $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its 5(5)555(5)5 ( 5 )-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d)-(e).
@@
(5.2.3). Let $n_{4}(v)=4$. By Lemma 2.26(f), $v$ has at most one 5(5)555(5)5 ( 5 )-neighbour. First, suppose that $v$ has no 5(5)555(5)5 ( 5 )-neighbours, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its other neighbours by R7(d)-(e). We may further assume that $v$ has exactly one 5(5)555(5)5 ( 5 )-neighbour, say $v_{1}$. Denote by $v_{2}$ and $v_{7}$ the common neighbours of $v$ and $v_{1}$. Recall that each of $v_{2},v_{7}$ is a $5^{+}$-vertex by Lemmas 2.17 and 2.19(a). Moreover, for $i\in\{2,7\}$, if $v_{i}$ is adjacent to a $4$-vertex, then $v_{i}$ is different from $5(4)$-vertex by Lemma 2.23(b). Recall also that a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex by Lemma 2.19(b).
@@
Suppose first that $v$ has a support vertex $x$. If $v$ has no $5(4)$-neighbour, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). If $v$ has a $5(4)$-neighbour, then $v$ has also a $7$-neighbour by Lemma 2.26(g). Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its 5(5)555(5)5 ( 5 )-neighbour by R7(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d).

### v3

@@ 2.2. The case $\Delta=7$
(2.1). Let $m_{3}(v)=1$. If $m_{4}(v)\leq 1$, then $v$ is incident to two $5^{+}$-faces, and by R2, $v$ receives $\frac{1}{5}$ from each of those $5^{+}$-faces. Thus, $\mu^{*}(v)\geq 2\times\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=2$, then $v$ has either a $7(5^{-})$-neighbour or two $7(6)$-neighbours by Lemma 2.22(a). In both cases, $v$ receives at least $\frac{1}{4}=\min\{\frac{1}{4},2\times\frac{1}{6}\}$ from its $7$-neighbours by R6-R8. On the other hand, $v$ receives $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq\frac{1}{4}+\frac{1}{5}-\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1. If $m_{4}(v)=3$, then $v$ has two $7(6^{-})$-neighbours by Lemma 2.22(b). By applying R6-R8, $v$ receives totally at least $\frac{1}{3}$ from its $7$-neighbours. Thus, $\mu^{*}(v)\geq\frac{1}{3}-\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to its incident $3$-face by R1.
@@
(3.2). Let $m_{3}(v)=4$. Suppose first that $m_{4}(v)=0$. By Lemma 2.25(d), $v$ has two neighbours different from $5^{-}$- and $6(6)$-vertices, i.e., $v$ has two neighbours consisting of $6(5^{-})$- or $7$-vertices. It follows that $v$ receives totally at least $\frac{1}{6}=\min\{2\times\frac{1}{9},\frac{1}{9}+\frac{1}{12},2\times\frac{1}{12}\}$ from its $6^{+}$-neighbours by R3-R9, and $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus, $\mu^{*}(v)\geq 1+\frac{1}{6}+\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Next we suppose that $m_{4}(v)=1$. By Lemma 2.25(e), $v$ has at least three $6^{+}$-neighbours different from $6(6)$-vertex. If $v$ has no $7(7)$-neighbour, then $v$ receives totally at least $3\times\frac{1}{9}$ from its $6^{+}$-neighbours by R3-R8. If $v$ has a $7(7)$-neighbours, then $v$ has four $6^{+}$-neighbours different from $6(6)$-vertex by Lemma 2.25(f), and so $v$ receives totally at least $4\times\frac{1}{12}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{1}{3}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(3.3.1). Let $n_{5}(v)=0$. By Lemma 2.26(b), $v$ has at most one $6(6)$-neighbour, i.e., $v$ has at least four $6^{+}$-neighbours different from $6(6)$-vertex. First, suppose that $v$ has a $6(6)$-neighbour. If $v$ has no $6(5^{-})$-neighbour, then all the neighbours of $v$ different from $6(6)$-vertex are $7$-vertices, and by applying R6-R9, $v$ receives totally at least $4\times\frac{1}{6}$ from its $7$-neighbours. So, $\mu^{*}(v)\geq 1+4\times\frac{1}{6}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has a $6(5^{-})$-neighbour, then, by Lemma 2.26(d), $v$ has three $7$-neighbours, at least one of which is a $7(5^{-})$-vertex. Then $v$ receives at least $\frac{1}{9}$ from its $6(5^{-})$-neighbour by R3-R5, at least $\frac{2}{9}$ from each of its $7(5^{-})$-neighbours by R6-R7, at least $\frac{1}{6}$ from each of its other $7$-neighbours by R8-R9. So, $v$ receives totally at least $\frac{1}{9}+\frac{2}{9}+2\times\frac{1}{6}=\frac{2}{3}$ from its $6^{+}$-neighbours. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Suppose now that $v$ has no $6(6)$-neighbour. That is, all neighbours of $v$ are $6^{+}$-vertices different from $6(6)$-vertex. By Lemma 2.26(b), $v$ has at least one $7$-neighbour, say $x$. Note that if $v$ has exactly four $6$-neighbours, then $x$ must be a $7(5^{-})$-vertex by Lemma 2.26(c). This means that $v$ has either four $6$-neighbours and one $7(5^{-})$-neighbour or at most three $6$-neighbours and two $7$-neighbours. In each case, $v$ receives totally at least $\frac{2}{3}=\min\{4\times\frac{1}{9}+\frac{2}{9},3\times\frac{1}{9}+2\times\frac{1}{6}\}$ from its $6^{+}$-neighbours by R3-R9. Thus, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
(3.3.2). Let $n_{5}(v)=1$. Notice that $v$ has no $6(6)$-neighbours by Lemma 2.26(e), i.e., all neighbours of $v$ but one are $7$- or $6(5^{-})$-vertices. Moreover, $v$ has at most two $6(5^{-})$-neighbours by Lemma 2.26(e). We distinguish three cases according to the number of $6(5^{-})$-vertices adjacent to $v$ as follows.
@@
Suppose next that $v$ has exactly one $6(5^{-})$-neighbour. Then $v$ has three $7$-neighbours, and by Lemma 2.26(f), one of which is a $7(5^{-})$-vertex. Thus, $v$ receives totally at least $\frac{1}{9}+2\times\frac{1}{6}+\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R9. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
Finally, suppose that $v$ has exactly two $6(5^{-})$-neighbours. Then, by Lemma 2.26(g), $v$ has two $7(5^{-})$-neighbours. So, $v$ receives totally at least $2\times\frac{1}{9}+2\times\frac{2}{9}$ from its $6^{+}$-neighbours by R3-R7. Hence, $\mu^{*}(v)\geq 1+\frac{2}{3}-5\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(5.1). Let $m_{3}(v)=4$. By Lemma 2.28(b), $v$ has at most six $4$-neighbours. If $n_{4}(v)\leq 4$, then $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), and at most $\frac{2}{9}$ to each of its other neighbours by R7. Consequently, $\mu^{*}(v)\geq 3-4\times\frac{1}{4}-3\times\frac{2}{9}-4\times\frac{1}{3}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1. Suppose now that $5\leq n_{4}(v)\leq 6$. Clearly, $v$ has no $5(5)$-neighbour, since a $4$ vertex is not adjacent to any $5(5)$-vertex by Lemma 2.21(a). If $v$ has a $3$-neighbour $x$, then the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.19. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Thus $\mu^{*}(v)\geq 3+\frac{1}{5}-6\times\frac{1}{4}-\frac{1}{5}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{5}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1. On the other hand, if $v$ has no $3$-neighbour, then $\mu^{*}(v)\geq 3-6\times\frac{1}{4}-\frac{1}{9}-4\times\frac{1}{3}>0$ after $v$ sends $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its neighbours other than $4$-vertex by R7, and $\frac{1}{3}$ to each of its incident $3$-faces by R1.
@@
(5.2). Let $m_{3}(v)=5$. Notice that $v$ has at most one $3$-neighbour, since a $3$-vertex does not incident to any $3$-faces by Lemma 2.19. First we assume that $v$ has a $3$-neighbour $x$. In such a case, all $3$-faces incident to $v$ must have a consecutive ordering. Moreover, the edge $xv$ is contained in at least one $5^{+}$-face by Lemma 2.19. So, $v$ receives at least $\frac{1}{5}$ from its incident $5^{+}$-face by R2. Observe first that if $v$ has three $4$-neighbours, then at least one of them must be a support neighbour of $v$. In such a case, $v$ has at most four poor neighbours by Lemma 2.28(h), and recall that a poor vertex receives at most $\frac{1}{4}$ from $v$ by R7(b),(c). It then follows that $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-4\times\frac{1}{4}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its poor neighbours by R7(b),(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e). We now suppose that $v$ has at most two $4$-neighbours. It then follows from Lemma 2.28(c) that $v$ has at most four poor neighbours. Thus, similarly as above, we have $\mu^{*}(v)\geq 3+\frac{1}{5}-5\times\frac{1}{3}-\frac{1}{5}-2\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{5}$ to its $3$-neighbour $x$ by R7(a), at most $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), and at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
@@
We may further assume that $v$ has no $3$-neighbours. By Lemma 2.28(b), $v$ has at most six $4$-neighbours, i.e., $n_{4}(v)\leq 6$. On the other hand, recall that $m_{3}(v)=5$. Therefore, if $v$ has more than four $4$-neighbours, then there exists $v_{i}\in N(v)$ such that $vv_{i}$ is contained in two $3$-faces, i.e., $v_{i}$ is a support neighbour.
@@
(5.2.1). Let $n_{4}(v)\leq 2$. Note that $v$ has at most four poor neighbours by Lemma 2.28(c). Then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-2\times\frac{1}{4}-2\times\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
@@
(5.2.2). Let $n_{4}(v)=3$. By Lemma 2.28(e), $v$ has at most two $5(5)$-neighbours. Suppose first that $v$ has exactly two $5(5)$-neighbours $x,y$. Since a $4$-vertex has no $5(5)$-neighbours by Lemma 2.21(a), we deduce that $x$ and $y$ are adjacent. On the other hand, a $5(5)$-vertex has at most one bad $5$-neighbour by Lemma 2.26(b). This implies that the neighbours of $v$ other than $4$- and $5(5)$-vertices are $5(3^{-})$- or $6^{+}$-vertices. Thus we have $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-2\times\frac{2}{9}-2\times\frac{1}{18}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). Suppose now that $v$ has at most one $5(5)$-neighbour. In such a case, we have again $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-3\times\frac{1}{4}-\frac{2}{9}-3\times\frac{1}{9}>0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to each of its $5(5)$-neighbours by R7(c), at most $\frac{1}{9}$ to each of its other neighbours by R7(d),(e).
@@
(5.2.3). Let $n_{4}(v)=4$. By Lemma 2.28(f), $v$ has at most one $5(5)$-neighbour. First, suppose that $v$ has no $5(5)$-neighbours, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-3\times\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), at most $\frac{1}{9}$ to each of its other neighbours by R7(d)-(e). We may further assume that $v$ has exactly one $5(5)$-neighbour, say $v_{1}$. Recall that each of $v_{2},v_{7}$ is a $5^{+}$-vertex by Lemmas 2.19 and 2.21(a). Moreover, for $i\in\{2,7\}$, if $v_{i}$ is adjacent to a $4$-vertex, then $v_{i}$ is different from $5(4)$-vertex by Lemma 2.25(b). Recall also that a $4(1^{+})$-vertex is adjacent to at most one $4$-vertex by Lemma 2.21(b).
@@
Suppose first that $v$ has a support neighbour $x$. If $v$ has no $5(4)$-neighbour, then $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-2\times\frac{1}{18}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), at most $\frac{1}{18}$ to each of its other neighbours by R7(e). If $v$ has a $5(4)$-neighbour, then $v$ has also a $7$-neighbour by Lemma 2.28(g). Thus, $\mu^{*}(v)\geq 3-5\times\frac{1}{3}-4\times\frac{1}{4}-\frac{2}{9}-\frac{1}{9}=0$ after $v$ sends $\frac{1}{3}$ to each of its incident $3$-faces by R1, $\frac{1}{4}$ to each of its bad $4$-neighbours by R7(b), $\frac{2}{9}$ to its $5(5)$-neighbour by R7(c), $\frac{1}{9}$ to its $5(4)$-neighbour by R7(d).